OpenStudy (anonymous):

If you have a question ask me. I cannot Guaranteed that i will know the answer but I will try my best.

OpenStudy (anonymous):

how would this look graphed? 2(x - y) 4 - 2y

OpenStudy (anonymous):

do you want a visual?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

ok. ill try

OpenStudy (anonymous):

What are the discontinuities of the function f(x) = the quantity x squared minus 16 over the quantity 6x minus 24 ?

OpenStudy (anonymous):

wait by the way, is this two equations?

OpenStudy (anonymous):

this is what it tells me Select the graph that best describes the following inequalities and system of inequalities. 2(x - y) 4 - 2y

OpenStudy (anonymous):

is there pictures with it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you post them it would be easier than drawing it all out. plz. and sorry

OpenStudy (anonymous):

nico are you wanting your question answered?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

does this help

OpenStudy (anonymous):

yes. now one more question hon, is the equation you gave me supposed to be 2 equations or one and you have to solve it?

OpenStudy (anonymous):

i have to solve it but don't know how and i think its just one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

megan do you know precal?

OpenStudy (anonymous):

i just figured something i think you need a < > in it? do you

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

um some. jacky

OpenStudy (anonymous):

ur a sophmore what classes are you taking

OpenStudy (anonymous):

math classes* im a sophomore too :)

OpenStudy (anonymous):

ok. shoot

OpenStudy (anonymous):

um im taking geometry right now but we just got into trig

OpenStudy (anonymous):

im getting off in 5 minutes btw

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but you can try me. and ill try

OpenStudy (anonymous):

\[\sin(x+\frac{ \pi }{ 6 })-\cos(x+\frac{ \pi }{ 3 }) =\sqrt{3}sinx\]

OpenStudy (anonymous):

maria i think your answer is the 4th one you sent. please dont hold it against me if its wrong

OpenStudy (anonymous):

its okay but thank you!!!!

OpenStudy (anonymous):

the square root of 3

OpenStudy (anonymous):

your welcome hon

OpenStudy (anonymous):

jacky did you hear me?

OpenStudy (anonymous):

By the sine and cosine subtraction formula: (a) sin(x - π/6) = sin(x)cos(π/6) - cos(x)sin(π/6) = sin(x)(√3/2) - cos(x)(1/2) = (√3/2)sin(x) - (1/2)cos(x). (b) cos(π/3 - x) = cos(π/3)cos(x) + sin(π/3)sin(x) = (1/2)cos(x) + (√3/2)sin(x). So, we have: [sin(x - π/6) + cos(π/3 - x)]/sin(x) = [(√3/2)sin(x) - (1/2)cos(x) + (1/2)cos(x) + (√3/2)sin(x)]/sin(x), from above = [√3*sin(x)]/sin(x), by simplifying = √3, by canceling sin(x). i hope this is correct.

OpenStudy (anonymous):

i have to go jackie but i hope this helps. if its not the answer your looking for im deeply sorry.

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

is one of your options x-4